# 3D Strain Field Reconstruction by Inversion of Dynamical Scattering

### v1

@@ Abstract
However, so far 3D strain fields have remained largely inaccessible to state-of-the-art strain measurement techniques in (scanning-) transmission electron microscopy (S/TEM). Several techniques exist like nano-beam electron diffraction (Usuda et al., 2005; Béché et al., 2009; Mahr et al., 2021), dark-field electron holography (Hÿtch et al., 2008; Béché et al., 2011), peak fitting (Bierwolf et al., 1993; Rosenauer et al., 1996), geometric phase analysis (Hÿtch et al., 2003), or other analyses of atomic resolution images (Niermann et al., 2011). Critically, all these techniques assume a strain field that is constant along the electron beam direction. Most of these techniques can only measure the strain components within the lateral plane, perpendicular to the electron beam, while the components in beam direction remain inaccessible. Furthermore, strain measurement techniques at atomic resolution often require thin specimen of a few ten nanometers at maximum, which alters the strain within the specimen due to surface relaxation effects (Jacob et al., 1998; Beyer et al., 2021).
For strain measurements the specimen must be oriented into strongly diffracting orientations, such as systematic row or zone-axis conditions. Under these orientations dynamical diffraction is dominant in samples thick enough to be representative of a bulk structure. For strain fields varying in beam direction dynamical diffraction results in a highly non-linear dependence of the measured signals on not only the strain inhomogeneity and its depth, but also on parameters like specimen thickness and the beam tilt (Lubk et al., 2014; Javon et al., 2014; Meißner et al., 2019), further complicating strain analysis with most techniques.
@@
$\frac{\partial}{\partial z}\phi_{g}(z;x,k_{\perp})=\mathrm{i}2\pi\left[s_{g}+g\epsilon(x,z)\right]\phi_{g}(z;x,k_{\perp})+\\ \mathrm{i}\frac{\pi}{k_{0}}\sum_{g}U_{g-g^{\prime}}\phi_{g^{\prime}}(z;x,k_{\perp}),$ (1)
@@
Fig. 4(b) shows the $qx$-plot obtained from this measurement. The experimental semiconvergence angle was 5.0 mrad, which corresponds to a probe size of roughly 0.4 nm. Clearly visible is the change within the $qx$-plots for $x$-coordinates between -25 nm and 25 nm, which is caused from the scattering at the inclined layer. Furthermore, a slight shifting of the reflections itself towards negative $q$-coordinates with increasing $x$-coordinates can be seen within the measurement. This is a measurement artifact caused by a misalignment of the beam shift-tilt decoupling within the microscope. Furthermore, a slight shift of the intensity patterns within the reflections to negative $q$ coordinates with increasing $x$-coordinates can be noticed. This is caused by a slight bending of the specimen (Niermann et al., 2025). Both artifacts have been corrected by incorporating an additional shear transformation within the $qx$-plane into the reconstruction. The lamella thickness $t=52.5\text{\nobreakspace nm}$ and the residual beam tilt were evaluated in advance by minimizing the loss function with respect to these parameters in the unstrained areas.
@@ Acknowledgements.
L. N. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project 492463633. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. We thank the Center for Electron Microscopy (ZELMI), Technische Universität Berlin, and the Structure Research and Electron Microscopy group, Humboldt-Universität zu Berlin, in the context of the Alliance Center Electron Microscopy (ACEM) for support. The ACEM is funded under the Excellence Strategy of the Federal Government and the Länder by the Berlin University Alliance (BUA).

### v2

@@ Abstract
However, so far 3D strain fields have remained largely inaccessible to state-of-the-art strain measurement techniques in (scanning-) transmission electron microscopy (S/TEM). Several techniques exist like nano-beam electron diffraction (Usuda et al., 2005; Béché et al., 2009; Mahr et al., 2021), dark-field electron holography (Hÿtch et al., 2008; Béché et al., 2011), peak fitting (Bierwolf et al., 1993; Rosenauer et al., 1996), geometric phase analysis (Hÿtch et al., 2003), or other analyses of atomic resolution images (Niermann et al., 2011). Critically, all these techniques assume a strain field that is constant along the electron beam direction. Most of these techniques can only measure the strain components within the lateral plane, perpendicular to the electron beam, while the components in beam direction remain inaccessible for a measurement along a single direction. Furthermore, strain measurement techniques at atomic resolution often require thin specimen of a few ten nanometers at maximum, which alters the strain within the specimen due to surface relaxation effects (Jacob et al., 1998; Beyer et al., 2021).
For strain measurements the specimen must be oriented into strongly diffracting orientations, such as systematic row or zone-axis conditions. Under these orientations dynamical diffraction is dominant in samples thick enough to be representative of a bulk structure. For strain fields varying in beam direction dynamical diffraction results in a highly non-linear dependence of the measured signals on not only the strain inhomogeneity and its depth, but also on parameters like specimen thickness and the beam tilt (Lubk et al., 2014; Javon et al., 2014; Meißner et al., 2019), further complicating strain analysis with most techniques. In contrast to electrons, X-rays do not exhibit strong dynamical diffraction effects, which enables advanced computational techniques like Bragg coherent diffraction imaging to be used for 3D strain measurements (Pfeifer et al., 2006; Cherukara et al., 2018).
@@
$\frac{\partial}{\partial z}\phi_{g}(z;x,k_{\perp})=\mathrm{i}2\pi\left[s_{g}+g\epsilon(x,z)\right]\phi_{g}(z;x,k_{\perp})+\\ \mathrm{i}\frac{\pi}{k_{0}}\sum_{g^{\prime}}U_{g-g^{\prime}}\phi_{g^{\prime}}(z;x,k_{\perp}),$ (1)
@@
Fig. 4(b) shows the $qx$-plot obtained from this measurement. The experimental semiconvergence angle was 5.0 mrad, which corresponds to a probe size of roughly 0.4 nm. Clearly visible is the change within the $qx$-plots for $x$-coordinates between -25 nm and 25 nm, which is caused from the scattering at the inclined layer. Furthermore, a slight shifting of the reflections itself towards negative $q$-coordinates with increasing $x$-coordinates can be seen within the measurement. This is a measurement artifact caused by a misalignment of the beam shift-tilt decoupling within the microscope. Furthermore, a slight shift of the intensity patterns within the reflections to negative $q$ coordinates with increasing $x$-coordinates can be noticed. This is caused by a slight bending of the specimen (Niermann et al., 2025). Both artifacts have been corrected by incorporating an additional shear transformation within the $qx$-plane into the reconstruction. The lamella thickness $t=52.5\text{\penalty 10000\ nm}$ and the residual beam tilt were evaluated in advance by minimizing the loss function with respect to these parameters in the unstrained areas.
@@ Acknowledgements.
L. N. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project 492463633. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. We thank the Center for Electron Microscopy (ZELMI), Technische Universität Berlin, and the Structure Research and Electron Microscopy group, Humboldt-Universit"at zu Berlin, in the context of the Alliance Center Electron Microscopy (ACEM) for support. The ACEM is funded under the Excellence Strategy of the Federal Government and the Länder by the Berlin University Alliance (BUA).